The 3rd harmonic frequency of a guitar string is 768 Hz.

a) Calculate its 1st harmonic (fundamental) frequency. (2 marks)

b) Calculate its 4th harmonic frequency. (2 marks)


1
Expert's answer
2021-01-29T05:19:56-0500

a)


"f_n=n\sdot\frac{v}{2L}=nf_1\to f_3=3f_1\to f_1=f_3/3=768/3=256(Hz)" . Answer


b)


"f_4=4f_1=4\cdot256=1024(Hz)" . Answer